OpenStudy (anonymous):

How to integrate the following: 2x/(x^3-1)

OpenStudy (anonymous):

\[I=\int\limits \frac{2 x }{\left( x-1 \right)\left( x^2+x+1 \right) }dx\] \[\frac{ 2x }{\left( x-1 \right)\left( x^2+x+1 \right) }=\frac{ A }{ x-1 }+\frac{ Bx+C }{ x^2+x+1 }\] \[2x=A \left( x^2+x+1 \right)+\left( Bx+C \right)\left( x-1 \right)\] equate terms like powers of x 0=A+B 2=A-B+C 0=A-C find A,B,C and integrate.

OpenStudy (anonymous):

can you complete it?

OpenStudy (anonymous):

http://prntscr.com/8an7mo this term is much complicated

OpenStudy (anonymous):

Thanks Surjithayer. Your reply was very helpful and I could complete it.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Unable to get correct Answer. Can you further solve To final answer.

OpenStudy (anonymous):

A+B=0,B=-A A-C=0 C=A 2=A+A+A \[A=\frac{ 2 }{ 3 },B=-A=-\frac{ 2 }{ 3 },C=A=\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

\[I=\int\limits \frac{ \frac{ 2 }{ 3 } }{ x-1 }dx+\int\limits \frac{ \frac{ -2 }{ 3 }x+\frac{ 2 }{ 3 } }{ x^2+x+1 }dx\] \[=\frac{ 2 }{ 3 }\ln \left( x-1 \right)-\frac{ 1 }{ 3 }\int\limits \frac{ 2x-2 }{ x^2+x+1 }dx\] \[=\frac{ 2 }{ 3 }\ln \left( x-1 \right)-\frac{ 1 }{ 3 }\int\limits \frac{ 2x+1-3 }{ x^2+x+1 }dx\] \[=\frac{ 2 }{ 3 }\ln \left( x-1 \right)-\frac{ 1 }{ 3 }\int\limits \frac{ 2x+1 }{ x^2+x+1 }dx+\int\limits \frac{ 1 }{x^2+x+1 }dx\]

OpenStudy (anonymous):

Thanks Surjithayer now it is clear.

OpenStudy (anonymous):

yw